OpenStudy (ajanijones):

Medals!!!! There are 10 marbles in a bag. Tyrell draws one marble out at a time without looking and replaces it in the bag. Tyrell did this 100 times. Tyrell drew red marbles 9 times, green marbles 53 times, and blue marbles 38 times. Based on this Tyrell's experiment, predict how many of each color marble is in the bag. A. 3 red, 3 green, and 4 blue B. 1 red,1 green, and 4 blue C. 3 red, 4 green, and 3 blue D. 1 red, 5 green, and 4 blue

OpenStudy (ajanijones):

@OliverSt4rz @OrangeMaster @Awesome781 @johnweldon1993 @Kamizamurai @SolomonZelman @Samuel91798 HELP PLZZZ..

OpenStudy (anonymous):

its no be because when you add up the tree numbers it equals six when there is 10 marbles in the bag

OpenStudy (anonymous):

and its not A because red was picked least so it would have the lease amount in bag. Does any of this help?

OpenStudy (anonymous):

50/100 10/100 40/100 Those were about his original results from picking from the bag 100 times. So, half are green, a tenth are red, and 4 tenths are blue. 5/10= 5 green 1/10= 1 red 4/10= 4 blue So, D is the answer.